Circular No. 7144

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 H1 Steven Lee reports his visual discovery of a comet at a star party near

Mudgee, New South Wales. The following observations are available:

1999 UT  α_{2000}  δ_{2000}  m_{1} Observer
Apr. 16.5  14.^{h}0  −69^{o}  9 Lee
16.59549 13 49^{m}24.^{s}64 −68 21′33′′.6  Garradd
16.59650 13 49 23.58 −68 21 33.3  _{”}
16.59941 13 49 20.80 −68 21 32.3  _{”}
16.61689 13 49 03.53 −68 21 26.4  _{”}
16.61777 13 49 02.67 −68 21 26.0  _{”}
16.61839 13 49 02.14 −68 21 25.7  _{”}

S. Lee (Mudgee, N.S.W.). Comet diffuse with no visible tail. G. J. Garradd (Loomberah). 0.45-m f/5.4 reflector + CCD. m_{2} = 13.9–

14.2. Visually, the comet has a 3′ coma slightly elongated northward.

HZ HERCULIS B. Margon, E. W. Deutsch, Z. M. Leinhardt, and S. F. Anderson,

University of Washington, report: “It has been noted (IAUC 7139) that the expected x-ray flux from Her X-1 has not been observed during the most recent 35-day precession cycle. Optical spectra of HZ Her obtained with the ARC 3.5-m telescope at Apache Point on Apr. 12 UT (system binary phase 0.41, where 0.5 is maximum light) show prominent C III/N III (464-, 465-nm) and He II (468.6-nm) emission lines, at strength comparable to that reported previously in the literature (e.g., Anderson et al. 1994, Ap.J. 436, 319). As the great strength of the C III/N III complex is known to be due to Bowen pumping triggered by intense He II (30.4 nm), and the 468.6-nm line has very high excitation potential, the conclusion seems inescapable that HZ Her is still irradiated by the x-ray source.”

SUPERNOVAE 1998bu, 1999D, 1999E, 1999X, 1999Z, 1999ac, 1999ai

CCD V magnitudes by L. Kiss and B. Csák, Piszkéstető, Hungary: SN 1998bu in NGC 3368, Mar. 19.922 UT, 18.3 ± 0.4; SN 1999D in NGC 3690, Mar. 19.863, 17.0 ± 0.3; SN 1999E, Mar. 19.955, 16.8 ± 0.3; SN 1999X in CGCG 180-22, Mar. 18.835, 18.3 ± 0.4; SN 1999Z in UGC 5608, Mar. 19.910, 17.3 ± 0.3; SN 1999ac in NGC 6063, Mar. 19.988, 14.2 ± 0.1; SN 1999ai, Mar. 19.939, 18.9 ± 0.4.

1999 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green